OpenStudy (anonymous):

compute: [\lim_{x\to 0}\frac{\sqrt{\cos(x)}-\cos(x)}{x^2}]

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{\sqrt{\cos(x)}-\cos(x)}{x^2}\]??

OpenStudy (anonymous):

yup, that one :) i just learned how to write it :)

OpenStudy (anonymous):

do you know L'hopital rule?

OpenStudy (anonymous):

just apply.

OpenStudy (anonymous):

we're not supposed to use it

OpenStudy (anonymous):

if it is have that form again, re apply

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

yeah i have seen these before "don't use l'hopital" why? who knows

OpenStudy (anonymous):

satellite73 here, sure you get

OpenStudy (anonymous):

i guess you can use power series, but that is even more complicated

OpenStudy (anonymous):

@satellite73 because we don't study it yet.

OpenStudy (anonymous):

lol, because we haven't learned it yet just like HOA said

OpenStudy (anonymous):

maybe you are supposed to do some trick like \[\frac{\sqrt{\cos(x)}}{x^2}-\frac{\cos(x)}{x^2}\]

OpenStudy (anonymous):

we need to use "squeeze" rule chain rule etc.

OpenStudy (anonymous):

strongly agree

OpenStudy (anonymous):

tried it but i can't seem to get rid of the \[x^{2}\]

OpenStudy (anonymous):

no my idea is bad

OpenStudy (anonymous):

i tried the "squeeze" principle but with no luck

OpenStudy (anonymous):

does it help to know that the answer is \(\frac{1}{4}\)?

OpenStudy (anonymous):

lol, no... I found the answer using wolfram but I need the whole way

OpenStudy (anonymous):

no idea?

OpenStudy (anonymous):

I am trying now. but no result yet. do you want to give up? if so, I have no reason to continue but if you try all your best, I am with you

OpenStudy (anonymous):

no I don't want to give up, it's math :)

OpenStudy (anonymous):

i am not good as others, I have just one thing: stubborn!!!hihi... ok, let's try

OpenStudy (anonymous):

thank you, I tried to multiply by \[\frac{ \sqrt{cosx}+cosx }{ \sqrt{cosx}+cosx } \] and got to \[\frac{ cosx - \cos^2x }{ x^2*(\sqrt{cosx}+cosx) }\] other way was to start from: \[-1<\cos x<\] and try to find a squeeze

OpenStudy (anonymous):

I dont' think squeeze works, but if you factor the numerator = cosx (1-cos x) then you can break the x^2 from denominator into 2 , first term is cosx/x * (1-cosx)x * (sqr....) .

OpenStudy (anonymous):

do you know the formula lim sinx/x =1 and lim (cos x -1)/x =0?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok, let try, you go on your way, i am on mine. if we stuck, you must wait for other's help.

OpenStudy (anonymous):

ok, thank you

OpenStudy (anonymous):

i've been trying to solve this allday, I think i'm missing some crucial trigonometric identity

OpenStudy (anonymous):

see if you can make some sense of this squeeze i tried \[-1 \le cosx \le 1\] \[-1 \le -cosx \le 1\] \[\sqrt{cosx}-1 \le \sqrt{cosx}-cosx \le \sqrt{cosx}+1\] \[\frac{\sqrt{cosx}-1}{x^2} \le \frac{\sqrt{cosx}-cosx}{x^2} \le \frac{\sqrt{cosx}+1}{x^2}\]

OpenStudy (anonymous):

sorry friend, cannot help

OpenStudy (anonymous):

@eidinovm I accidentally wrote the result here... http://openstudy.com/study#/updates/5154969be4b0507ceba113be

OpenStudy (anonymous):

\[ L=\lim_{x\to0}\sqrt{\cos x}\frac{1-\sqrt{\cos x}}{x^2}\times\frac{1+\sqrt{\cos x}}{1+\sqrt{\cos x}}\\ L=\lim_{x\to0}\sqrt{\cos x}\frac{1-\cos{x}}{x^2(1+\sqrt{\cos x})}\times\frac{1+\cos x}{1+\cos x}\\ L=\lim_{x\to0}\sqrt{\cos x}\frac{1-\cos^2x}{x^2(1+\sqrt{\cos x})(1+\cos x)}\\ L=\lim_{x\to0}\sqrt{\cos x}\times\left(\frac{\sin x}{x}\right)^2\times{1\over(1+\sqrt{\cos x})(1+\cos x)}\\ L=\sqrt{\cos 0}\times(1)^2\times{1\over (1+\sqrt{\cos 0})(1+\cos 0)}\\ \boxed{\Huge L={1\over 4}} \]

OpenStudy (anonymous):

thank you very much!!!!!